Lenter Supply Company is a medium-sized distributor of wholesale hardware supplies in the central Ohio area. It has been a client of yours for several years and has instituted excellent internal controls for sales at your recommendation.

In providing control over shipments, the client has prenumbered “warehouse removal slips” that are used for every sale. It is company policy never to remove goods from the warehouse without an authorized warehouse removal slip. After shipment, two copies of the warehouse removal slip are sent to billing for the computerized preparation of a sales invoice. One copy is stapled to the duplicate copy of a prenumbered sales invoice, and the other copy is filed numerically. In some cases, more than one warehouse removal slip is used for billing one sales invoice. The smallest warehouse removal slip number for the year is 11741 and the largest is 34687. The smallest sales invoice number is 45302 and the largest is 65747.

In the audit of sales, one of the major concerns is the effectiveness of the controls in ensuring that all shipments are billed. You have decided to use audit sampling in testing internal controls.

Required

  1. State an effective audit procedure for testing whether shipments have been billed. What is the sampling unit for the audit procedure?

  2. Assume that you expect no exceptions in the sample but are willing to accept a TER of 4 percent. At a 5 percent ARO, what is the appropriate sample size for the audit test? You may complete this requirement using attributes sampling.

  3. Design a random selection plan for selecting the sample from the population, using either systematic sampling or computer generation of random numbers. Use the sample size determined in part b. If you use systematic sampling, use a random starting point of 11932.

  4. Your supervisor suggests the possibility of performing other sales tests with the same sample as a means of efficiently using your audit time. List two other audit procedures that can conveniently be performed using the same sample, and state the purpose of each of the procedures.

  5. Is it desirable to test the occurrence of sales with the random sample you have designed in part c.? Why or why not?
Answer #1

The appropriate audit procedure for testing whether shipments have been billed is:

A sample of warehouse removal slip should be selected. It should be examined to see whether it has the proper sales invoice attach or not. So the sampling unit for the audit procedure will be warehouse removal slip.

Using attribute sampling method if auditor is willing to accept TER of 4% and at ARO of 5% without exception in the sampling size the sample size would be 75

Design of random selection plan from population:

Population size of ware house removal slips =34,687-11,741=22,946

Sample size determined in b = 75

Interval = ( Population sized)/(Sample sized determined)= ( 22,946)/75 =305

Consider the random starting point from 11878 warehouse of removal slip then random sample will be 111932, 12237, 12542, 12847, 13152

Other audit procedures that can be performed conveniently using same sample are as follows:

1.Test of accuracy: All attached sales invoices should be tested for checking clerical errors and accuracy.

2.Test of timing: Delay in time should be tested between warehouse removal slip date and billing date of timeline of billing.

Testing the occurrence of sales with the random sample we have designed in part C is not desirable this is because of following reason:

1.Auditor already knows that inventory was shipped in for the sales.

2.For checking the sales, the sales invoice entries in the sales journal is the sampling unit. As warehouse removal slips are not identical with sales invoices numbers it will not be desirable to use same sample